find the missing value for this problem it shows (6g___)(6g+3h)=36g^2-9h^2
Answer by jorel555(1290)

In general;x^2-y^2=(x+y)(x-y).
In this instance:
sqrt 36g^2=6g
sqrt 9h^2=3h
Thus
36g^2-9h^2=(6g+3h)(6g-3h)!!!!!!!!!